Grade 11Mechanics

What is the average velocity and average acceleration in the interval AC when the particle covers a semicircle (i.e if AC is a diameter) at the same speed?

Kumar Ankur
8 Years agoGrade 11

1 Answer

Eshan
8 Years ago
Let the speed bev.

Hence the average velocity =\dfrac{Displacement}{Time}=\dfrac{2R}{\pi R/v}=\dfrac{2v}{\pi}

Average acceleration=\dfrac{Change in velocity}{Time}=\dfrac{v-(-v)}{t}=\dfrac{2v}{\pi R/v}=\dfrac{2v^2}{\pi R}

=\dfrac{4v^2}{\pi (AC)}